If the perimeter of a tennis court is 228 feet and the length is 6 feet longer than twice the width, then what are the length and width?
Answer #1
x=width=36 2x+6=length=78 ---------------- p=2w+2L 2x+2(2x+6)=228 2x+4x+12=228 6x=216 x=36
answered by: Anderson
Answer #2

x=width=36 2x+6=length=78 ---------------- p=2w+2L 2x+2(2x+6)=228 2x+4x+12=228 6x=216 x=36